OpenStudy (anonymous):

What's the limit as x approaches infinity, sqrt(x^4-2x^3-2)/(10x^2+9)

OpenStudy (jamesj):

Divide top and bottom by x^2, then ...

OpenStudy (anonymous):

Since the top and bottom both go to infinity, you can use l'Hopital's rule! \[\lim_{x \rightarrow c} \frac{f(x)}{g(x)} = \lim_{x \rightarrow c} \frac{f'(x)}{g'(x)}\]

OpenStudy (jamesj):

In this case, it's liking shooting a fly with the gun of l'Hopital's rule. Not necessary. Follow my procedure and you see that the term becomes sqrt(1-2/x-2/x^4)/(10+9/x^2) Now what is the limit of that as x -> infinity?

OpenStudy (anonymous):

1/10?

OpenStudy (anonymous):

You can't divide the numerator by \(x^2\) because it's inside a square root.

OpenStudy (anonymous):

oh ok...

OpenStudy (jamesj):

@Buck: yes, exactly @Dalvoron: yes you can, notice that in doing so the term in the sqrt is divided by x^4.

OpenStudy (anonymous):

Touche.

OpenStudy (jamesj):

1/x^2 . sqrt(x^4-2x^3-2) = sqrt(1/x^4 . (x^4-2x^3-2) ) = ....

OpenStudy (anonymous):

so what if there wasn't a denominator? would it still apply?